OpenStudy (anonymous):

If an object is dropped from a height of 144 feet, the function h(t)= -16t^2+144 gives the height of the object after t seconds. When will the object hit the ground? A. 9 s B. 6 s C. 1.5 s D. 3 s

OpenStudy (whpalmer4):

well, solve the equation for h = 0...what positive value of t makes it so?

OpenStudy (whpalmer4):

at t = 0, h(0) = -16(0)^2 + 144 = 144 feet, which is the height from which it is being dropped. at some later t, h(t) = 0, and that's when it hits the ground.

OpenStudy (whpalmer4):

\[h(t) = 0 = -16t^2+144\]\[16t^2=144\]\[t=\]

OpenStudy (anonymous):

so it 9

OpenStudy (whpalmer4):

does 16*9*9 = 144?

OpenStudy (whpalmer4):

always, always check your answers in the original formula to make sure they correctly answer the question.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

it is 3 gj

OpenStudy (whpalmer4):

16(3)(3) = 144, so that is a reasonable answer. Note that t = -3 also satisfies the equation, but does not satisfy the conditions described in the problem, so it is not a correct solution here.

OpenStudy (anonymous):

ok